# CheMixHub: Datasets and Benchmarks for Chemical Mixture Property Prediction

### v1

@@ 3.4 Dataset splitting strategies
  • Temperature splits: As highlighted in Table 1, multiple tasks in CheMixHub have a temperature dependency. It is also desirable for industry to be able to predict the properties of certain mixtures in different temperature ranges than the training one. We bin the temperature range in 5 categories based on the temperature distribution observed across the 11 tasks and use the bins as a 5-fold split.
@@ 4.1 Modeling space
We benchmark three common embedding techniques: GNNs, CLMs and molecular descriptors. For GNNs, we use a GraphNets architecture [4] trained end-to-end with the other module to learn the embeddings during the task (details in Section A.2), while for CLMs we rely on frozen pre-trained representations from MolT5 [18]. All training is performed using the Adam optimizer [30]. Cheminformatics molecular descriptors are normalized 200 dimensional from RDKit obtained from descriptastorus [29]. After obtaining the molecular embeddings, we consider three different ways of infusing the molecular context into them: 1) element-wise multiplication, 2) concatenation, 3) feature-wise linear modulation (FiLM) layer [42].
@@
We explore two possible interaction modules: DeepSets [72] and self-attention [60], which can be thought of as enabling one-body and two-body interactions between molecules in the mixture, respectively. We consider six different types of permutation invariant aggregation operations: mean, max, attention-based aggregation [60], principal neighborhood aggregation (PNA) [16], PNA scaled according to the number of component in the mixture and set2set [63] which yields the mixture embedding. We then consider two different ways of incorporating the mixture context into it: 1) concatenation 2) using a FiLM layer.
@@
We consider using either a fully-connected predictive head that directly outputs a predicted value, or for tasks that consists in predicting properties which are temperature dependent, predicting the parameters of the Arrhenius equation which is then used to determine the final outputted value (see Section 4.4).
@@
To establish a strong non-deep learning baseline, we provide comparison with a gradient-boosted random forest model, XGBoost [12] using RDKit descriptors or MolT5 embeddings molecular features. These features are linearly combined with their respective molecular context and then concatenated with the overall mixture context to form the input for XGBoost (details in Section A.6).
@@ 4.2 Performances across tasks
Molecular rep. Mixture rep. Miscible Solvents Drug Solubility SPE NIST-full
$\rho$ $\Delta H_{mix}$ $\Delta H_{vap}$ $\ln(S)$ $\ln(\kappa)$ $\ln(\eta)$
@@
Molecular rep. Mixture rep. IlThermo MON NIST Olfaction
$\ln(\kappa)$ $\ln(\eta)$ MON $\ln(\eta)$ Mixture similarity
@@
We observe that traditional tree-based methods like XGBoost are robust baselines on a great variety of tasks. It is interesting to note that XGBoost-based methods greatly struggle on predicting $\Delta H_{mix}$ and $\Delta H_{vap}$, two properties well known for their non-linear mixture behavior. Overall, we observe pre-trained representations of CLMs like MolT5 tend to yield better performances across our dataset compared to GNN-based representation and cheminformatics descriptors. We believe that pre-training on related data would greatly improve the performances of GNNs, as it has been shown to in the literature [58, 54]. Regarding the choice of mixture interaction module, the need for higher level of interactions tend to be task-dependent, with no consistent advantage of one method over the other.
@@ 4.3 Generalization to new mixture sizes and molecules
4.4 Explicit physics-based modeling
Previous work have reported that incorporating known physical or chemical constraints into ML models can improve accuracy and generalizability of model predictions [7, 80]. We investigate how swapping a regular fully connected predictive head for a “physics-based” predictive head impacts the model’s performance. Namely, we modify the predictive head to output the coefficients of a physics law suitable for the task. For the scope of this paper, we limit our study to temperature-dependent tasks whose target property can be effectively modeled by the Arrhenius equation.
@@ 4.4 Explicit physics-based modeling
We observe that adding a physics bias via the Arrhenius equation greatly improves the performance of deep learning architectures in this setting. This technique also allows better interpretability of the predictive model, as it grounds it in known equations. Additionally, we note that the performance of XGBoost, a model that typically performs well on randomized splits, has significantly decreased performance when evaluated on a harder and more realistic split, highlighting the importance of generalizability assessment and not relying on a single split toi report in literature.
@@ 5 Conclusion
In this study, we introduced CheMixHub, a comprehensive suite of datasets and benchmarks designed to accelerate research in chemical mixture property prediction. Addressing the critical need for standardized resources in a field characterized by scattered datasets, inconsistent evaluation protocols, and limited open-source model implementations, CheMixHub provides a curated collection of 11 tasks, diverse splitting strategies for robust generalization assessment, and initial baselines using representative ML models. Our work aims to lower the barrier to entry and foster systematic progress in understanding and modeling these complex multi-molecular systems.
Our benchmarking revealed that traditional models like XGBoost with appropriate chemical features offer strong baselines on random splits, often rivaling more complex deep learning methods. This highlights the necessity for deep learning approaches to demonstrate clear advantages, particularly on more challenging out-of-distribution tasks. We observed that datasets with greater monomolecular diversity (e.g., MON, Olfactory Similarity) benefit from hierarchical modeling as well as tasks with well known non-linear relationships, such as enthalpy of mixing $\Delta H_{mix}$ and heat of vaporization $\Delta H_{mix}$, underscoring the need for advanced modeling and rigorous evaluation beyond simple random splits. Encouragingly, the explicit incorporation of physics-based constraints, like the Arrhenius equation for temperature-dependent properties, significantly enhanced model performance and generalization, suggesting a fruitful direction for future work in fusing domain knowledge with data-driven techniques. The optimal level of interaction modeling—whether one-body (DeepSets-like) or explicit many-body approaches—also remains task-dependent and warrants further investigation, alongside innovations in aggregation, context conditioning, and attention mechanisms tailored for mixtures.

### v2

@@ 3.4 Dataset splitting strategies
  • Temperature splits: As highlighted in Table 1, multiple tasks in CheMixHub have a temperature dependency. It is also desirable for industry to be able to predict the properties of certain mixtures in different temperature ranges than the training ones. We bin the temperature range into 5 categories based on the temperature distribution observed across the 11 tasks and use the bins as a 5-fold split.
@@ 4.1 Modeling space
We benchmark three common embedding techniques: GNNs, CLMs and molecular descriptors. For GNNs, we use a GraphNets architecture [battaglia2018relational] trained end-to-end with the other module to learn the embeddings during the task (details in Section A.4), while for CLMs we rely on frozen pre-trained representations from MolT5 [edwards2022translation]. All training is performed using the Adam optimizer [kingma2014adam]. Cheminformatics molecular descriptors are normalized 200 dimensional from RDKit obtained from descriptastorus [Kelley2024-kd]. After obtaining the molecular embeddings, we consider three different ways of infusing the molecular context into them: 1) element-wise multiplication, 2) concatenation, 3) feature-wise linear modulation (FiLM) layer [perez2018film].
@@
We explore two possible interaction modules: DeepSets [zaheer2017deep] and self-attention [vaswani2017attention], which can be thought of as enabling one-body and two-body interactions between molecules in the mixture, respectively. We consider six different types of permutation invariant aggregation operations: mean, max, attention-based aggregation [vaswani2017attention], principal neighborhood aggregation (PNA) [corso2020principal], PNA scaled according to the number of component in the mixture and set2set [vinyals2015order] which yields the mixture embedding. We then consider two different ways of incorporating the mixture context into it: 1) concatenation 2) using a FiLM layer.
@@
We consider using either a fully-connected predictive head that directly outputs a predicted value, or for tasks that are known to be modeled by an Arrhenius relationship, predicting the parameters of the Arrhenius equation and using it to determine the final outputted value (see Section 4.4).
@@
To establish a strong non-deep learning baseline, we provide comparison with a gradient-boosted random forest model, XGBoost [chen2016xgboost] using RDKit descriptors or MolT5 embeddings molecular features. These features are linearly combined with their respective molecular context and then concatenated with the overall mixture context to form the input for XGBoost (details in Section A.8).
@@ 4.2 Performances across tasks
Molecular
rep.
Mixture
rep.
Miscible Solvents Drug Solubility SPE NIST-full
$\rho$ $\Delta H_{\mathrm{mix}}$ $\Delta H_{\mathrm{vap}}$ $\ln(S)$ $\ln(\kappa)$ $\ln(\eta)$
@@
Molecular
rep.
Mixture
rep.
IlThermo Fuel mixtures NIST Olfactory mixtures
$\ln(\kappa)$ $\ln(\eta)$ MON $\ln(\eta)$ Perceptual similarity
@@
We observe that traditional tree-based methods like XGBoost are robust baselines on a great variety of tasks. It is interesting to note that XGBoost-based methods greatly struggle on predicting $\Delta H_{\mathrm{mix}}$ and $\Delta H_{\mathrm{vap}}$, two properties well known for their non-linear mixture behavior. Overall, we observe pre-trained representations of CLMs like MolT5 tend to yield better performances across our dataset compared to GNN-based representation and cheminformatics descriptors. We believe that pre-training on related data would greatly improve the performances of GNNs, as it has been shown to in the literature [tom2025molecules, sharma2023formulation]. Regarding the choice of mixture interaction module, the need for higher level of interactions tend to be task-dependent, with no consistent advantage of one method over the other.
@@ 4.3 Generalization to new mixture sizes and molecules
4.4 Explicit physics-based modeling improves performances
Previous work have reported that incorporating known physical or chemical constraints into ML models can improve accuracy and generalizability of model predictions [bradford2023chemistry, zhu2024differentiable]. We investigate how swapping a regular fully connected predictive head for a “physics-based” predictive head impacts the model’s performance. Namely, we modify the predictive head to output the coefficients of a physics law suitable for the task. For the scope of this paper, we limit our study to temperature-dependent tasks whose target property can be effectively modeled by the Arrhenius equation.
@@ 4.4 Explicit physics-based modeling improves performances
We observe that adding a physics bias via the Arrhenius equation greatly improves the performance of deep learning architectures in this setting. This technique also allows better interpretability of the predictive model, as it grounds it in known equations. Additionally, we note that the performance of XGBoost, a model that typically performs well on randomized splits, significantly decreases when evaluated on a harder and more realistic split, highlighting the importance of generalizability assessment and not relying on a single split to report in literature.
4.5 Transfer learning capabilities of models within the Miscible Solvent dataset tasks.
We evaluate transfer learning capabilities of the best performing models on each of the 3 property prediction tasks (Density $\rho$, $\Delta H_{\mathrm{mix}}$ and $\Delta H_{\mathrm{vap}}$) of the Miscible Solvents dataset. For each task, we finetune the best models of the other two tasks and compare their performance to the original best model for this task reported in Section 4.2.
Table 4: Intra-dataset transfer learning capabilities depend on task difficulty Metrics are reported on 5-fold random CV splits. The mean and standard deviation are reported. The original best model statistics are taken from Section 4.2 and Appendix A.9.
Fine-tuning Dataset
Best model
Original Dataset
Pearson $\rho$ ($\uparrow$) MAE ($\downarrow$) Kendall $\tau$ ($\uparrow$)
$\rho$ $\rho$ 0.999 ± 0.000 0.003 ± 0.000 0.973 ± 0.000
$\Delta H_{\mathrm{mix}}$ 0.955 ± 0.006 0.021 ± 0.001 0.824 ± 0.009
$\Delta H_{\mathrm{vap}}$ 0.929 ± 0.008 0.026 ± 0.002 0.769 ± 0.018
$\Delta H_{\mathrm{vap}}$ $\Delta H_{\mathrm{vap}}$ 0.999 ± 0.000 0.071 ± 0.002 0.976 ± 0.001
$\Delta H_{\mathrm{mix}}$ 0.808 ± 0.017 1.063 ± 0.057 0.611 ± 0.034
$\rho$ 0.644 ± 0.088 1.366 ± 0.176 0.465 ± 0.074
$\Delta H_{\mathrm{mix}}$ $\Delta H_{\mathrm{mix}}$ 0.976 ± 0.003 0.157 ± 0.002 0.835 ± 0.002
$\Delta H_{\mathrm{vap}}$ 0.626 ± 0.022 0.527 ± 0.008 0.439 ± 0.025
$\rho$ 0.348 ± 0.044 0.629 ± 0.013 0.237 ± 0.033
We observe dramatic differences depending on the task the model was originally trained on: Models initially trained to predict highly non-linear properties — harder tasks — like $\Delta H_{\mathrm{mix}}$ and $\Delta H_{\mathrm{vap}}$ perform really well when finetuned to predict density $\rho$ but the model initially trained to predict density $\rho$ fails at delivering good performances on $\Delta H_{\mathrm{mix}}$ and $\Delta H_{\mathrm{vap}}$ predictions. Architectural differences may also play a role in this phenomenon. We perform additional inter-dataset transfer learning experiments in Appendix A.10.
4.6 Zero-shot capabilities across viscosity prediction tasks
We performed additional experiments to investigate the zero-shot capabilities of the best performing deep learning models for each of the viscosity ($\ln(\eta)$) prediction tasks in CheMixHub and observe good zero shot capabilities for tasks that have similar viscosity value ranges.
Table 5: Zero shot learning capabilities of models across the Viscosity $\mathbf{\ln(\eta)}$ prediction tasks in CheMixHub Metrics are reported on 5-fold random CV splits. The mean and standard deviation are reported. The original best model statistics are taken from Section 4.2 and Appendix A.9.
Zero-shot Dataset
Best model
Original Dataset
Pearson $\rho$ ($\uparrow$) MAE ($\downarrow$) Kendall $\tau$ ($\uparrow$)
NIST NIST 0.991 ± 0.001 0.030 ± 0.001 0.939 ± 0.001
NIST-full 0.985 ± 0.002 6.806 ± 0.012 0.926 ± 0.004
IlThermo $\ln(\eta)$ 0.575 ± 0.028 5.880 ± 0.129 0.451 ± 0.024
NIST-full NIST-full 0.992 ± 0.000 0.055 ± 0.000 0.966 ± 0.000
IlThermo $\ln(\eta)$ 0.775 ± 0.018 0.811 ± 0.078 1.000 ± 0.000
NIST 0.694 ± 0.021 6.281 ± 0.005 1.000 ± 0.000
IlThermo $\ln(\eta)$ IlThermo $\ln(\eta)$ 0.995 ± 0.001 0.076 ± 0.002 0.968 ± 0.001
NIST-full 0.956 ± 0.004 0.276 ± 0.032 0.880 ± 0.015
NIST 0.452 ± 0.041 4.815 ± 0.030 0.330 ± 0.047
@@ 5 Conclusion
In this study, we introduced CheMixHub, a comprehensive suite of datasets and benchmarks designed to accelerate research in chemical mixture property prediction. Addressing the critical need for standardized resources in a field characterized by scattered datasets, inconsistent evaluation protocols, and limited open-source model implementations, CheMixHub provides a curated collection of 11 tasks, diverse splitting strategies for robust generalization assessment, and initial baselines using representative ML models. Our work aims to lower the barrier to entry and foster systematic progress in understanding and modeling these complex multi-molecular systems. Our benchmarking revealed that traditional models like XGBoost with appropriate chemical features offer strong baselines on random splits, often rivaling more complex deep learning methods. This highlights the necessity for deep learning approaches to demonstrate clear advantages, particularly on more challenging out-of-distribution tasks. We observed that datasets with greater monomolecular diversity (e.g., fuel and olfactory mixtures) benefit from hierarchical modeling as well as tasks with well known non-linear relationships, such as enthalpy of mixing $\Delta H_{\mathrm{mix}}$ and heat of vaporization $\Delta H_{\mathrm{mix}}$, underscoring the need for advanced modeling and rigorous evaluation beyond simple random splits. Encouragingly, the explicit incorporation of physics-based constraints, like the Arrhenius equation for temperature-dependent properties, significantly enhanced model performance and generalization, suggesting a fruitful direction for future work in fusing domain knowledge with data-driven techniques. The optimal level of interaction modeling—whether one-body (DeepSets-like) or explicit many-body approaches—also remains task-dependent and warrants further investigation, alongside innovations in aggregation, context conditioning, and attention mechanisms tailored for mixtures.
@@ 6 Limitations
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We clearly state contributions in the introduction and answers the research questions introduced there.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Limitations of our work are discussed in the conclusion section.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: The paper does not include theoretical results.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Scripts to process data and run models are provided on GitHub and detailed in the paper and appendix.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: All scripts and results are provided on the project’s GitHub. If the dataset could not be directly included on the GitHub, clear instructions of where to find the data and how to set up the data processing pipleine are provided.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: Training and test details are described in the main body of the paper. The hyperparameter search space is described in Appendix. The splits are provided on GitHub for reproducibility.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: All results provided in the benchmark are conducted on multiple splits of the data, and confidence intervals are reported.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: Compute resources details described in the Appendix
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The authros have reviewed the NeurIPS Code of Ethics and consider this work is conform to it.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Positive and negative impacts of this work have been discussed throughout the introduction and conclusion sections.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: The models explored are not high risk for misuse.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: The code to process data is open source. For each dataset, the original paper is cited in the main publication as well as on the GitHub and the license name is provided in the supplementary material.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: For new dataset, code for generation is provided and the process is described in Appendix. For the new benchmark, all results are provided on GitHub.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A] .
  69. Justification: the paper does not involve crowdsourcing nor research with human subjects
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A] .
  74. Justification: The paper does not involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer: [N/A]
  79. Justification: LLMs are not used in this paper. We only explore the use of language-based molecular representation.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.